### v1

@@ E.1 The Algorithm
where the last implication follows from the fact that $(1-\frac{s+1}{p^{t}})\geq\frac{1}{2}\cdot\left(\frac{20}{19}\right)^{2}$ for all $s$, since $t=\lceil\log_{p}(20s^{2})\rceil+1$. We have set $\tau\leq\frac{\epsilon^{2}}{40p^{t}}$ in the algorithm. Hence, the number of heavy buckets is at least $(s+1)$.
Appendix F Proof of Theorem F.8
In this section we will prove the generalized version of Theorem 1.3 and Theorem 1.4 for $\mathcal{G}=\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}$, where $p_{1},\ldots,p_{t}$ are distinct primes. Let $\omega_{p_{i}}$ be a primitive $p_{i}^{th}$ root of unity, $i\in[t]$. For any multivariate polynomial $\sum_{i_{1}=0}^{d_{1}}\cdots\sum_{i_{t}=0}^{d_{t}}g_{i_{1},\ldots,i_{t}}x_{1}^{i_{1}}\cdots x_{t}^{i_{t}}=g(x_{1},\ldots,x_{t})\in\mathbb{C}[x_{1},\ldots,x_{t}]$, let $|g|_{1}\;:=\;\sum_{i_{1}=0}^{d_{1}}\cdots\sum_{i_{t}=0}^{d_{t}}|g_{i_{1},\ldots,i_{t}}|$.
The following definition helps us to understand what we mean by $\mu$-close to $(m_{1},\ldots,m_{t})$-granular.
Definition F.1.
($\mu$-close to $(m_{1},\ldots,m_{t})$-granular) A complex number $v$ is said to be $\mu$-close to $(m_{1},\ldots,m_{t})$-granular with respect to $\mathbb{Z}_{p_{1}}\times\cdots\times\mathbb{Z}_{p_{t}}$ if $|v-\frac{g(\omega_{p_{1}},\ldots,\omega_{p_{t}})}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}|\leq\mu$, for some $g(X_{1},\ldots,X_{t})\in\mathbb{Z}[X_{1},\ldots,X_{t}]$, which is a non-zero multivariate polynomial of degree at most $(p_{1}-1)\cdots(p_{t}-1)$, and $\omega_{p_{i}}$ is a $p_{i}^{th}$ root of unity, $i\in[t]$.
Now let us state the generalized form of Theorem 1.3 for $\mathcal{G}$.
Theorem F.2.
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function. Then, if $f$ is $\mu$-close to $s$-sparse in $\ell_{2}$, for each character $\chi\in B$, $\widehat{f}(\chi)$ is $\frac{\mu}{\sqrt{s}}$-close to $(\lceil\log_{p_{1}}(s)^{\frac{1}{t}}\rceil+1,\ldots,\lceil\log_{p_{t}}(s)^{\frac{1}{t}}\rceil+1)$-granular.
The following theorem is a corollary of Theorem F.2 which gives us a lower bound on the Fourier coefficients of $f:\mathcal{G}\to\{-1,+1\}$.
Theorem F.3 (Theorem F.8, restated).
Let $p_{1},\ldots,p_{t}$ be distinct primes, and let $f:\mathbb{Z}_{p_{1}}^{n_{1}}\times\mathbb{Z}_{p_{2}}^{n_{2}}\cdots\times\mathbb{Z}_{p_{t}}^{n_{t}}\to\{-1,+1\}$. Then, for each $\chi\in\operatorname{supp}(f)$, we have
$\left|\widehat{f}(\chi)\right|\geq\frac{1}{(m^{2}s_{f})^{\lceil\varphi(m)/2\rceil}},$
where $m:=p_{1}\dots p_{t}\text{ and }\varphi(m)=(p_{1}-1)(p_{2}-1)\cdots(p_{t}-1)$ and $s_{f}$ is the sparsity of the function $f$.
The following theorem is the generalized form of Theorem 1.4 for $\mathcal{G}$.
Theorem F.4.
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function such that $f$ is $\mu\leq\frac{1}{8(m^{2}s)^{\varphi(m)}}$ close to $s$-sparse in $\ell_{2}$ norm, where $m=p_{1}\cdots p_{t}$ and $\varphi(m)=(p_{1}-1)\cdots(p_{t}-1)$. Then, there exists an $s$-sparse Boolean valued function $F:\mathcal{G}\to\{-1,+1\}$ which is $\sqrt{2}\mu$ close to $f$ in $\ell_{2}$.
To prove Theorem F.2, we need the following lemma.
Lemma F.5.
$P_{r_{1}+H_{1},\ldots,r_{t}+H_{t}}f(x_{1},\ldots,x_{t})=\frac{1}{|H_{1}^{\perp}\cdots H_{t}^{\perp}|}\sum_{(z_{1},\ldots,z_{t})\in H_{1}^{\perp}\cdots H_{t}^{\perp}}f(x_{1}-z_{1},\ldots,x_{t}-z_{t})\chi((z_{1},\cdots,z_{t})),$
where $\chi$ is a character of $\mathcal{G}$ and for each $i$, $H_{i}$ is a subspace (hence a subgroup) of $\mathbb{Z}_{p_{i}}^{n_{i}}$ and $x_{i},z_{i}\in\mathbb{Z}_{p_{i}}^{n_{i}}$.
The following lemma plays a crucial role in proving Lemma F.5.
Claim 4 (Restated Lemma A.10).
$\sum_{h\in H}\chi_{h}=|H|\cdot 1_{H^{\perp}}$, where $H=H_{1}\times\cdots\times H_{t}$ is a subgroup of $\mathcal{G}$, and $h=(h_{1},\ldots,h_{t}),\ h_{i}\in H_{i}\ \forall i$. That is,
$\sum_{h_{1}\in H_{1},\ldots,h_{t}\in H_{t}}\chi_{h_{1}}\cdots\chi_{h_{t}}=|H_{1}|\cdots|H_{t}|\cdot 1_{H_{1}^{\perp}}\cdots 1_{H_{t}^{\perp}}.$
Here, for each $i$, $H_{i}$ is a subgroup of $\mathbb{Z}_{p_{i}}^{n_{i}}$ and hence a subspace $\mathbb{Z}_{p_{i}}^{n_{i}}$, as $\mathbb{Z}_{p_{i}}^{n_{i}}$ is a vector space.
Proof F.6 (Proof of Lemma F.5).
$P_{r_{1}+H_{1},\ldots,r_{t}+H_{t}}f(x_{1},\ldots,x_{t})$
$=\sum_{\alpha_{1}\in r_{1}+H_{1},\ldots,\alpha_{t}\in r_{t}+H_{t}}\hat{f}(\alpha_{1},\ldots,\alpha_{t})\chi^{(p_{1})}_{\alpha_{1}}(x_{1})\cdots\chi^{(p_{t})}_{\alpha_{t}}(x_{t})$
$=\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\hat{f}(r_{1}+\beta_{1},\ldots r_{t}+\beta_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}+\beta_{i}}(x_{i}),\text{ where }\alpha_{i}=r_{i}+\beta_{i},i\in\{1\ldots,t\}\text{ and }\beta_{i}\in H_{i}\forall i$
$=\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\hat{f}(r_{1}+\beta_{1},\ldots,r_{t}+\beta_{t})\prod_{i=1}^{t}\bigg{(}\chi^{(p_{i})}_{r_{i}}(x_{i})\chi^{(p_{i})}_{\beta_{i}}(x_{i})\bigg{)}$
$=\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\hat{f}(r_{1}+\beta_{1},\ldots,r_{t}+\beta_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i})$
$=\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i})\frac{1}{|\mathcal{G}|}\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\overline{\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}+\beta_{i}}(y_{i})}$
$=\frac{1}{|\mathcal{G}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i})\prod_{i=1}^{t}\overline{\chi^{(p_{i})}_{r_{i}+\beta_{i}}(y_{i})}$
$=\frac{1}{|\mathcal{G}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots,y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(-y_{i})\sum_{\beta_{1}\in H_{1},\ldots,\beta_{t}\in H_{t}}\prod_{i=1}^{t}\chi^{(p_{i})}_{\beta_{i}}(x_{i}-y_{i})$
$=\frac{1}{|\mathcal{G}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{y_{1}\in\mathbb{Z}_{p_{1}}^{n_{1}},\ldots y_{t}\in\mathbb{Z}_{p_{t}}^{n_{t}}}f(y_{1},\ldots,y_{t})\bigg{(}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(-y_{i})\bigg{)}|H_{1}|\cdots|H_{t}|1_{H_{1}^{\perp}}(x_{1}-y_{1})\cdots 1_{H_{t}^{\perp}}(x_{t}-y_{t})$
$=\frac{1}{|H_{1}^{\perp}|\cdots|H_{t}^{\perp}|}\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i})\sum_{x_{1}-y_{1}\in H_{1}^{\perp},\ldots x_{t}-y_{t}\in H_{t}^{\perp}}f(y_{1},\ldots,y_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(-y_{i})$
$=\frac{1}{|H_{1}^{\perp}|\cdots|H_{t}^{\perp}|}\sum_{x_{1}-y_{1}\in H_{1}^{\perp},\ldots,x_{t}-y_{t}\in H_{t}^{\perp}}f(y_{1},\ldots,y_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(x_{i}-y_{i})$
$=\frac{1}{|H_{1}^{\perp}|\cdots|H_{t}^{\perp}|}\sum_{z_{1}\in H_{1}^{\perp},\ldots z_{t}\in H_{t}^{\perp}}f(x_{1}-z_{1},\ldots,x_{t}-z_{t})\prod_{i=1}^{t}\chi^{(p_{i})}_{r_{i}}(z_{i}),$
where $x_{i}-y_{i}=z_{i}$ for all $i\in[t]$.
Now we prove Theorem F.2.
Proof F.7 (Proof of Theorem F.2).
Fix a $\chi^{(\mathcal{G})}:=\chi^{(p_{1})}\dots\chi^{(p_{t})}\in B$, where $\chi^{(p_{i})}$ is a character of $\mathbb{Z}^{n_{i}}_{p_{i}}$ for all $i\in[t]$, and $B$ is $B$ is the set of character functions corresponding to the $s$ largest coefficients. Set for each $i\in[t]$, $m_{i}:=\lceil\log_{p_{i}}(s)^{\frac{1}{t}}\rceil+1$.
Let $A_{i}^{\perp}\subset\mathbb{Z}^{n_{i}}_{p_{i}}$, be the set of solutions to $A_{i}\chi=0$, for all $i\in[t]$. Let $H_{i}$ be the coset of $A_{i}^{\perp}$ that are solutions to system of linear equations $A_{i}\chi=b_{i}$. From the definition of projection operator, we have
$P_{H_{1},\ldots,H_{t}}f(x_{1},\ldots,x_{t})=\sum_{\chi^{(p_{1})}\in H_{1},\ldots,\chi^{(p_{t})}\in H_{t}}\widehat{f}(\chi^{(p_{1})},\ldots,\chi^{(p_{t})})\chi^{(p_{1})}(x_{1})\cdots\chi^{(p_{t})}(x_{t}),$ (F.1)
where $\chi_{r_{j}}^{(p_{i})}$ is a character of $\mathbb{Z}_{p_{i}}^{n_{i}}$ (follows from Definition A.1).
We need to show that for each $\chi_{r_{j}}=(\chi_{r_{j}}^{(p_{1})}\cdots\chi_{r_{j}}^{(p_{t})})\in B$, where $\chi_{r_{j}}^{(p_{i})}$ is a character of $\mathbb{Z}_{p_{i}}^{n_{i}}$, there exists $m_{i}\times n_{i}$ matrices $A_{i}$ and a $m_{i}$-dimensional vector $b_{i}$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A_{1}\chi^{(p_{1})}\,=\,b_{1},\ldots,A_{t}\chi^{(p_{t})}\,=\,b_{t}\right\}=\{\chi_{r_{j}}\},$ (F.2)
and, the $\ell_{2}$ Fourier weight of $S\cap H$ is upper bounded by $\frac{\mu^{2}}{s}$, where $H=H_{1}\times\cdots\times H_{t}$, that is,
$\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\leq\frac{\mu^{2}}{s},$
where $B$ is the set of character functions corresponding to the $s$ largest coefficients and $S$ is the set of rest of the characters, and $H_{i}$ is the solution set of the system of linear equations $A_{i}\chi=b_{i}$ for all $i\in[t]$. Let $E_{1}$ and $E_{2}$ be the first and the second events respectively. So we need to prove that both these events occur simultaneously with non-zero probability.
To show the existence of $A_{i}$ and $b_{i}$ for all $i\in[t]$, given $\chi_{r_{j}}\in B$, for which Equation F.2 is true we will need to consider two cases:
Case 1: $\chi_{r_{j}}=0$
. Observe that for all $A_{i}$, $i\in[t]$, we have $A_{i}\chi_{r_{j}}^{(p_{i})}=0$. If entries of $A_{i}$ are sampled independently and uniformly at random from $\mathbb{Z}_{p_{i}}$, for each $i\in[t]$. Then, for all $\chi_{r_{q}}\in B\setminus\{\chi_{r_{j}}\}$, we have
$\Pr_{A_{1},\ldots,A_{t}}\left[A_{1}\chi_{r_{q}}^{(p_{1})}=0,\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}=0\right]=\frac{1}{p_{1}^{m_{1}}\cdots p_{t}^{m_{t}}}\leq\frac{1}{s}.$
Note that $\Pr_{A_{1},\ldots,A_{t}}\left[E_{1}\right]$ denotes the probability of event $E_{1}$ when $A_{i}$ are the random matrices generated by the above-mentioned procedure. Using union bound, we get
$\Pr_{A_{1},\ldots,A_{t}}\left[\exists q\neq j\text{ such that }A_{1}\chi_{r_{q}}^{(p_{1})}=0,\ldots,A_{t}\chi_{r_{q}}^{(p_{t})}=0\right]\leq\frac{s-1}{s}<1.$
Therefore, if $\chi_{r_{j}}^{(p_{i})}=0$ for all $i\in[t]$, then there exists $m_{i}\times n_{i}$ matrices $A_{i}$ such that

### v2

@@ 6.1 The Algorithm
where the last implication follows from the fact that $(1-\frac{s+1}{\Delta^{t}})\geq\frac{1}{2}\cdot\left(\frac{20}{19}\right)^{2}$ for all $s$, since $t=\lceil\log_{\Delta}(20s^{2})\rceil+1$. We have set $\tau\leq\frac{\epsilon^{2}}{40\Delta^{t}}$ in the algorithm. Hence, the number of heavy buckets is at least $(s+1)$.
7 Degree over $\mathcal{G}$
The following definition is a generalization of $\deg_{2}(f)$, that is degree over $\mathbb{Z}_{2}$ of a Boolean function $f$, as defined in [BC99].
Definition 7.1.
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function. Consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a coset of $\mathcal{G}$ as defined in Definition 2.17. Then the degree over $\mathcal{G}$ of $f$, denoted by $\deg_{\mathcal{G}}$, is defined in the following way.
$\deg_{\mathcal{G}}(f)=$ $\min_{t}\{\lfloor\log_{\Delta}|\mathcal{G}|\rfloor-t:\exists r_{1},\ldots,r_{t}\in\operatorname{supp}(\widehat{f}),b=(b_{1},\ldots b_{t})\in\mathbb{Z}_{\Delta}^{t},$
$\text{ such that }f|_{V_{b,r_{1},\ldots,r_{t}}}\text{ has full sparsity}\},$
where $\Delta=\operatorname{LCM}\{p_{1}^{m_{1}},\ldots,p_{T}^{m_{T}}\}$.
Remark 7.2.
In Definition 7.1, the restriction $f|_{V_{b,r_{1},\ldots,r_{t}}}$ has full sparsity means that $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}\geq\frac{|\mathcal{G}|}{\Delta^{t}}\geq\Delta^{\lfloor\log_{\Delta}|\mathcal{G}|\rfloor-t}$ from Lemma 2.22 and Lemma 2.19, where $s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}$ is the Fourier sparsity of the function $f|_{V_{b,r_{1},\ldots,r_{t}}}$.
Next, let us define the $p_{i}^{m_{i}}\times p_{i}^{m_{i}}$ matrices $V_{i}$, $i\in[T]$ by
$V_{i}=\begin{bmatrix}1&1&1&\cdots&1\\ 1&\omega_{p_{i}^{m_{i}}}&\omega_{p_{i}^{m_{i}}}^{2}&\cdots&\omega_{p_{i}^{m_{i}}}^{{p_{i}^{m_{i}}}-1}\\ 1&\omega_{p_{i}^{m_{i}}}^{2}&\omega_{p_{i}^{m_{i}}}^{4}&\cdots&\omega_{p_{i}^{m_{i}}}^{2({p_{i}^{m_{i}}}-1)}\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ \cdot&\cdot&\cdot&\cdot&\cdot\\ 1&\omega_{p_{i}^{m_{i}}}^{{p_{i}^{m_{i}}}-1}&\omega_{p_{i}^{m_{i}}}^{2({p_{i}^{m_{i}}}-1)}&\cdots&\omega_{p_{i}^{m_{i}}}^{({p_{i}^{m_{i}}}-1)({p_{i}^{m_{i}}}-1)}\\ \end{bmatrix}$
Let us consider the Kronecker product (Definition 2.37) of these matrices $V_{1},\ldots,V_{T}$, which is a $(p_{1}^{m_{1}}\cdots p_{T}^{m_{T}})\times(p_{1}^{m_{1}}\cdots p_{T}^{m_{T}})$ matrix, and is denoted by $V_{1}\otimes\cdots\otimes V_{T}$.
Lemma 7.3.
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function and $\widehat{f}$ denotes its Fourier transform. Then $\widehat{f}=\frac{1}{p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}}(V_{1}\otimes\cdots\otimes V_{T})f$, where $f$ is the $p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}\times 1$ matrix formed by the values $f(x)$, $x\in\mathcal{G}$ are arranged in usual lexicographic order, and $\widehat{f}$ is the $p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}\times 1$ matrix formed by the values $\widehat{f}(\chi_{r})$, $r\in\mathcal{G}$ are arranged in usual lexicographic order.
Proof.
We will use induction on $T$ to prove this.
Base case.
When $T=1$, then for each element $i\in\mathbb{Z}_{p_{1}^{m_{1}}}$, the Fourier coefficients are given by
$\widehat{f}(\chi_{i})=\frac{1}{p_{1}^{m_{1}}}\biggl{(}f(0)+f(1)\omega_{p_{1}^{m_{1}}}^{i}+f(2)\omega_{p_{1}^{m_{1}}}^{2i}+\ldots+f(p_{1}^{m_{1}}-1)\omega_{p_{1}^{m_{1}}}^{(p_{1}^{m_{1}}-1)i}\biggr{)}.$
Hence $\widehat{f}=\frac{1}{p_{1}^{m_{1}}}V_{1}f$ in this case.
Induction step.
Let for all $T\leq K-1$, $\widehat{f}=\frac{1}{p_{1}^{m_{1}}\cdots p_{K-1}^{m_{K-1}}}V_{1}\otimes\cdots\otimes V_{K-1}f$.
For $T=K$, for each element $(i,r)\in\mathbb{Z}_{{p_{1}}^{m_{1}}}\times\cdots\times\mathbb{Z}_{{p_{K-1}}^{m_{K-1}}}\times\mathbb{Z}_{{p_{K}}^{m_{K}}}$, the Fourier coefficients of $f$ are given by
$\widehat{f}(\chi_{(i,r)})=\frac{1}{p_{1}^{m_{1}}\cdots p_{K}^{m_{K}}}\sum_{x\in\mathbb{Z}_{{p_{1}}^{m_{1}}}\times\cdots\times\mathbb{Z}_{{p_{K-1}}^{m_{K-1}}}}$ $\biggl{(}\bigl{(}f(0,x)+f(1,x)\omega_{p_{K}^{m_{K}}}^{i}+f(2,x)\omega_{p_{K}^{m_{K}}}^{2i}+\cdots$
$+f({p_{K}^{m_{K}}}-1,x)\omega_{p_{K}^{m_{K}}}^{({p_{K}^{m_{K}}}-1)i}\bigr{)}\omega_{\operatorname{LCM}\{p_{1}^{m_{1}},\ldots,p_{K-1}^{m_{K-1}}\}}^{r*x}\biggr{)}.$ (7.4)
From Induction step., any block corresponding to the coefficients $\widehat{f}(\chi_{i,r})$ and the function values $f(j,x)$ is given by the matrix $\omega_{p_{K}^{m_{K}}}^{ij}V_{1}\otimes\cdots\otimes V_{K-1}$, $r,x\in\mathbb{Z}_{{p_{1}}^{m_{1}}}\times\cdots\times\mathbb{Z}_{{p_{K-1}}^{m_{K-1}}}$. Hence, $\widehat{f}=\frac{1}{p_{1}^{m_{1}}\cdots p_{K}^{m_{K}}}V_{1}\otimes\cdots\otimes V_{K}f$.
Thus by induction it follows that $\widehat{f}=\frac{1}{p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}}V_{1}\otimes\cdots\otimes V_{T}f$ for any positive integer $T$. ∎
One can equivalently define $\deg_{\mathcal{G}}$ in the following way.
Definition 7.5.
Let $f:\mathcal{G}\to\{-1,+1\}$ be a Boolean valued function. Let us arrange the elements $x\in\mathcal{G}$ in lexicographic order along the rows and columns of the matrix $V_{1}\otimes\cdots\otimes V_{T}$. Consider all possible restrictions $f|_{V_{b,r_{1},\ldots,r_{t}}}$ of $f$, where $V_{b,r_{1},\ldots,r_{t}}$ is a coset of $\mathcal{G}$ as defined in Definition 2.17. Then the degree over $\mathcal{G}$ of $f$ is defined as the dimension of the largest coset $V_{b,r_{1},\ldots,r_{t}}$ such that the restriction $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value 11 at those points such that each entry of the row/column matrix formed by the sum of the rows/columns corresponding to the elements where $f|_{V_{b,r_{1},\ldots,r_{t}}}$ takes the value 11 is $\neq 0$.
[Note that in that case all the Fourier coefficients are nonzero.]
Remark 7.6.
Observe, from the Vandermonde matrix $V_{1}\otimes\cdots\otimes V_{T}$, that, any character, (except $\chi_{0}$, the empty character), of $f$ can be written as $\frac{1}{p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}}\biggl{(}\sum_{i}c_{i}\omega_{\Delta}^{i}-\sum_{j}d_{j}\omega_{\Delta}^{j}\biggr{)}$, where $c_{i},d_{j}$ are non-negative integers for all $i,j$ and $\omega_{\Delta}$ is a primitive root of unity, $\Delta=\operatorname{LCM}\{p_{1}^{m_{1}},\ldots,p_{T}^{m_{T}}\}$. (Any root of unity of the form $\omega_{p_{1}^{m_{1}}}\cdots\omega_{p_{T}^{m_{T}}}$ is also a $\omega_{\Delta}^{th}$ root of unity.) Note that the sum of elements of each row/column of $V_{1}\otimes\cdots\otimes V_{T}$ is equal to 0. That is, $\sum_{i}c_{i}\omega_{\Delta}^{i}+\sum_{j}d_{j}\omega_{\Delta}^{j}=0$. So, for a column with $\sum_{i}c_{i}\omega_{\Delta}^{i}\neq 0$, we have the corresponding character $\frac{1}{p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}}\biggl{(}\sum_{i}c_{i}\omega_{n}^{i}-\sum_{j}d_{j}\omega_{\Delta}^{j}\biggr{)}=\frac{1}{p_{1}^{m_{1}}\cdots p_{T}^{m_{T}}}\biggl{(}\sum_{i}c_{i}\omega_{\Delta}^{i}+\sum_{i}c_{i}\omega_{\Delta}^{i}\biggr{)}\neq 0$. Hence Definition 7.5 follows.
Note that, the first row/column, corresponding to the constant coefficient, is not equal to 0. In the case when $\mathcal{G}$ is of prime order, the constant coefficient is always $\neq 0$, as $f$ can take the value 11 at an odd number of elements and the value 1-1 at an even number of elements, or vice versa. In case of general Abelian groups, this may not be true, so the restriction has to be chosen in such a way that it does not take the value 11 at half of the points, so that the constant coefficient is not equal to 0.
8 Query lower bound on adaptively testing sparsity
Theorem 8.1 (Theorem 1.9, restated).
For Boolean valued functions on $\mathcal{G}$, to adaptively test $s$-sparsity, the query lower bound of any algorithm is $\Omega(\sqrt{s})$.
Let $\tau>0$ and $C$ be a constant dependent on $\tau$, and will be determined later. Let $H$ be a random subgroup (see Definition 2.38) of $\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}$ of co-dimension $t$ (see Definition 2.20), where $i_{1},\ldots,i_{Ct}\in\{1,\ldots,T\}$, and $\mathcal{C}$ is the set of all cosets of $H$. That is, $(H,\mathcal{C})$ forms a random $t$-dimensional coset structure (see Definition 2.39). Now, let us define the following two probability distributions:
  • We choose a random function $f$ randomly from $\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}$, conditioned on the fact that $f$ is $2-\tau$ far in $\ell_{2}$ from any function which has $\deg_{\mathcal{G}}=t$. Let us call this distribution $\mathcal{D}_{No}$.
Theorem 8.2.
Any adaptive query algorithm that distinguishes $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ with probability $\geq\frac{1}{3}$ has to make at least $\Omega(\Delta^{\frac{t}{2}})$ queries.
The following lemma will be useful in proving Theorem 8.2.
Lemma 8.3.
For any Boolean valued function $f:\mathcal{G}\to\{-1,+1\}$,
$\Delta^{r(f)}\geq s_{f}\geq\Delta^{\deg_{\mathcal{G}}(f)}.$
Proof.
The proof of the first inequality follows trivially from the definitions of sparsity $s_{f}$ (Definition 2.11) and dimension $r(f)$ (Definition 2.16) of $f$.
The proof of the second inequality follows from the definition of $\deg_{\mathcal{G}}(f)$ (Definition 7.1) and Remark 7.2, since $s_{f}\geq s_{f|_{V_{b,r_{1},\ldots,r_{t}}}}\geq\Delta^{\lfloor\log_{\Delta}|\mathcal{G}|\rfloor-t}=\Delta^{\deg_{\mathcal{G}}(f)}$, where $V_{b,r_{1},\ldots,r_{t}}$ is the coset, with $t$ being the least possible nonnegative integer, such that $f|_{V_{b,r_{1},\ldots,r_{t}}}$ has full sparsity. ∎
Corollary 8.4 (Corollary of Lemma 8.3).
The set of functions with $r(f)=t$ is a subset of the set of functions with sparsity $s=\Delta^{t}$, which in turn is a subset of the set of functions with $\deg_{\mathcal{G}}=t$.
Theorem 1.9 directly follows from Theorem 8.2 and Lemma 8.3.
We will show that if an adaptive query algorithm makes less than $q<\Omega(\Delta^{t/2})$ queries, then the total variation distance $\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}$ between the two distributions $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ is $\leq\frac{1}{3}$, where the total variation distance is as defined in Definition 2.44 and Lemma 2.45. This proves that any adaptive query algorithm which distinguishes between $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$, that is, for which $\mathcal{D}_{Yes}-\mathcal{D}_{No}||_{TV}>\frac{1}{3}$ must make at least $\Omega(\Delta^{t/2})$ queries.
Any adaptive query algorithm that makes $q$ queries can be thought of as a $n^{\prime}$-ary decision tree, whose internal nodes are elements from $\mathbb{Z}_{p_{i_{1}}^{m_{i_{1}}}}\times\cdots\times\mathbb{Z}_{p_{i_{Ct}}^{m_{i_{Ct}}}}$ and the leaves are labeled by accept or reject, where $n^{\prime}=\operatorname{LCM}\{p_{i_{1}}^{m_{i_{1}}},\ldots,p_{i_{Ct}}^{m_{i_{Ct}}}\}$. Note that $n^{\prime}$ can be at most $\Delta$. Also, the length of each path from the root to the leaves of the tree is equal to $q$ and no element repeats in the tree, that is, it does not appear more than once in the tree.
Let $\mathcal{L}_{Yes}$ and $\mathcal{L}_{No}$ be the distributions on the leaves of the decision tree from $\mathcal{D}_{Yes}$ and $\mathcal{D}_{No}$ respectively. We will show that $\mathcal{L}_{Yes}-\mathcal{L}_{No}||_{TV}\leq\frac{1}{3}$ when the number of queries $q\leq\frac{1}{\sqrt{20}}\Delta^{\frac{t}{2}}$.
Lemma 8.5.
$\mathcal{L}_{Yes}-\mathcal{L}_{Uniform}||\leq\frac{1}{19}$.
Proof of Lemma 8.5.
Consider the set of all functions of dimension $t$. Let the depth $q$ of the decision tree corresponding to our algorithm is $\leq\frac{1}{\sqrt{20}}\Delta^{\frac{t}{2}}$. Then $t\geq 2\log_{\Delta}q+\log_{\Delta}20$. So by Theorem 2.36 (part 2), all the characters on a particular path fall into different buckets with probability $\geq(1-\frac{1}{20})$. Let us choose a function $f$ randomly from $\mathcal{D}_{Yes}$. For each leaf $\ell$, let us consider the characters on the path from root to $\ell$. Since the length of the path is $q$, so with probability $1-\frac{1}{20}$, the characters on the path belongs to different buckets. Then the probability that $f$ is consistent with the path to $\ell$, given that all the characters on the path from root to $\ell$ belong to different buckets is $\frac{1}{n^{\prime q}}$. Note that $n^{\prime}\leq n$. Therefore, the probability that $f$ is consistent with the path to $\ell$ is $\geq(1-\frac{1}{20})\frac{1}{n^{\prime q}}=\frac{19}{20}\frac{1}{n^{\prime q}}$. Let us denote this probability by $\mathcal{P}_{Yes}$.
Now, for each path from root to $\ell$, we know that $\mathcal{P}_{Uniform}(x)=\frac{1}{n^{\prime q}}$, where $x$ is a character in the path from root to $\ell$. Then,